LASER SPECTROSCOPY OF NiI: NEW ELECTRONIC STATES AND HYPERFINE STRUCTURE

\underline{A. S-C. CHEUNG}, H.F. PANG, W.S. TAM, J. YE, AND J. W-H. LEUNG, Department of Chemistry, The University of Hong Kong, Pokfulam Road, Hong Kong..

Two new electronic transition systems of NiI namely: the [14.0] ^{2}Φ_{7/2} - X^{2}Δ_{5/2} and the [15.7] ^{2}Δ_{5/2} - X^{2}Δ_{5/2} transitions were observed and analyzed using laser vaporization/reaction supersonic free jet expansion and high resolution laser induced fluorescence spectroscopy. In addition, the (v, 0) bands with v = 6 - 10 of the previously identified [14.6] ^{2}Δ_{5/2} - X^{2}Δ_{5/2} transition were found to be perturbed by the [15.7] ^{2}Δ_{5/2} state. All observed spectra show partially resolved hyperfine structure. Hyperfine width of rotational lines decreases rapidly as J increases suggested that the hyperfine structure for the [14.0] ^{2}Φ_{7/2}, the [14.6] ^{2}Δ_{5/2} and the [15.7] ^{2}Δ_{5/2} states conform to the Hund’s case a_{β} coupling scheme. The interaction between the [14.6]^{2}Δ_{5/2} and the [15.7]^{2}Δ_{5/2} states is evident in the progressive increase in hyperfine width in rotational lines of the [14.6] ^{2}Δ_{5/2} - X^{2}Δ_{5/2} transition as the vibrational quantum number increases. Deperturbation procedures were successfully applied to analyze the interaction between these two states. Accurate molecular and hyperfine constants for the [14.0] ^{2}Φ_{7/2}, the [14.6] ^{2}Δ_{5/2} and the [15.7] ^{2}Δ_{5/2} states were obtained and interpreted.